# Li and Zhu

### v1

@@
The surge in data availability has inundated decision-makers with an overwhelming array of choices. While existing approaches focus on optimizing decisions based on quantifiable metrics, practical decision-making often requires balancing measurable quantitative criteria with unmeasurable qualitative factors embedded in the broader context. In such cases, algorithms can generate high-quality recommendations, but the final decision rests with the human, who must weigh both dimensions. We define the process of selecting the optimal set of algorithmic recommendations in this context as human-centered decision making. To address this challenge, we introduce a novel framework called generative curation, which optimizes the true desirability of decision options by integrating both quantitative and qualitative aspects. Our framework uses a Gaussian process to model unknown qualitative factors and derives a diversity metric that balances quantitative optimality with qualitative diversity. This trade-off enables the generation of a manageable subset of diverse, near-optimal actions that are robust to unknown qualitative preferences. To operationalize this framework, we propose two implementation approaches: a generative neural network architecture that produces a distribution $\pi$ to efficiently sample a diverse set of near-optimal actions, and a sequential optimization method to iteratively generates solutions that can be easily incorporated into complex optimization formulations. We validate our approach with extensive datasets, demonstrating its effectiveness in enhancing decision-making processes across a range of complex environments, with significant implications for policy and management.
@@ 1 Introduction
With the explosion of data availability, human decision-makers faces an ever-growing set of information and decision choices. This phenomenon has spurred significant interest in utilizing data-driven machine learning (ML) methods to assist in human decision-making, and there is a quickly expanding body of literature dedicated to developing methods capable of recommending the optimal decision across a variety of scenarios based on a particular objective (e.g. Bertsimas and Kallus 2016, Kube et al. 2019, Bryan and Ventresca 2020).
@@
In this paper, we focus on developing decision recommendations while explicitly considering the context that these algorithmic suggestions are subject to human arbitration. We define this category of decision-making algorithms as human-centered decision making. As illustrated in Table 1, this category contrasts with common optimal decision making algorithms that seeks the best decision excluding any human agency (e.g. Zhao et al. 2012, Moodie et al. 2012, Zhu et al. 2017, Luedtke and van der Laan 2016, Athey and Wager 2021). We instead formulate a collaborative model where algorithms provide potential solutions that humans then evaluate, ensuring that algorithmic insights augment, rather than replace, human expertise, particularly in complex, nuanced decision scenarios.
@@
Algorithm gives decision input? No N/A Human Decision Making e.g. Doctors, Policymakers
Yes Optimal Decision Making e.g. optimal treatment rules Human-Centered Decision Making e.g. modeling to generate alternatives, generative curation
Current approaches in human-centered decision making often utilize ad-hoc constraints to heuristically generate a range of quantitatively near-optimal actions for decision-makers to choose from (Chang et al. 1982, Greistorfer et al. 2008). Alternatively, researchers try to elicit multiple dimensions of human preference and present different solutions along a Pareto frontier (Masin and Bukchin 2008, Lin et al. 2022, Wu et al. 2024).
We propose a different approach by recognizing that the ultimate goal of any recommendation system within a human-centered decision-making framework should be to facilitate decisions that are deemed most desirable by the human decision-maker themselves. Therefore, the goal of any set of $m$ recommendations $A_{1},\cdots,A_{m}$ should be to maximize the human desirability $\ell(\cdot)$ of the best recommendation
$\max_{A_{1},\cdots,A_{m}}\max_{i=1,\cdots,m}\ell(A_{i}).$ ($\mathcal{P}^{0}$)
If the human desirability $\ell(\cdot)$ is known, the problem directly degenerates to the setting of optimal decision making, where the goal is to find a single recommendation $A$ that maximizes $\ell(A)$.
However, human desirability of a decision is often not only affected by measurable, quantitative objectives, but also elusive qualitative factors from the broader problem context. We demonstrate this through several motivating examples:
Example 1.1 (Non-pharmaceutical Interventions for Pandemic Control)
During the COVID-19 pandemic, governments worldwide implemented a wide range of non-pharmaceutical interventions to manage the spread of the virus. Policymakers faced the challenge of balancing numerous quantitative factors, such as pandemic spread, control effectiveness, and fairness, with qualitative considerations like population acceptability, implementation feasibility, and political implications. Quantifying certain qualitative aspects is nearly impossible, and even for quantitative measures, policymakers often struggle to precisely determine their relative importance to formulate concrete objectives. Therefore, the implemented actions are often quantitatively suboptimal yet reflects a range of real-world desirability factors that are difficult to consider in model formulations. (Bertsimas et al. 2022)
Example 1.2 (Police Districting)
In police districting, departments typically organize their patrol forces by dividing a city’s geographical region into multiple patrol areas, known as zones. Quantitatively, the objective is to find an optimal districting plan that minimizes workload variance across zones (Larson 1974, Larson and Odoni 1981, Zhu et al. 2020, 2022, Xing et al. 2023). However, qualitative factors such as political and economic considerations and the shape of the districts are challenging to fully capture in a mathematically well-defined manner (Shirabe 2009). This complexity renders a comprehensive quantitative assessment of the underlying desirability of actions nearly impossible (Gardner et al. 2014). Figure 1(c) demonstrates a case where Atlanta police department rejected a quantitatively optimal plan due to ease of access to some highways and traffic constraints within the boundaries.
@@
(a) Pre-2019 plan
(b) Adopted plan
(c) Rejected plan
An example showing the difficult-to-define qualitative desirability through Atlanta police redistricting. Gray lines represent the basic geographical units patrolled by the police, the color depth indicates the police workload of each unit. Black lines outline the districting plans, and dashed red lines highlight the changes made to the pre-2019 plan in (a). The map in (b) is the new plan adopted by the APD after 2019. The map in (c) is quantitatively optimal (the workload variance across zones is minimized) according to an optimization model but was ultimately rejected by the APD because it overlooked traffic constraints and inadvertently cut off access to some highways with its zone boundaries.
Therefore, human desirability of any decision includes both explicitly known quantitative objectives and also unknown qualitative desirability factors. When we incorporate both known and unknown factors into $\ell(\cdot)$, Problem $\mathcal{P}^{0}$ is distinct from the classic optimal decision making problem.
To solve Problem $\mathcal{P}^{0}$, instead of directly maximizing over the set of recommended decisions – which becomes computationally prohibitive, especially when $m$ is large – we employ a distributional approach where we aim to create the optimal generative distribution $\pi$ to sample recommendations. In other words, the generative distribution $\pi$ should maximize the human desirability of the best action within a set of sampled recommendations sampled from $\pi$:
$\max_{\pi}~{}\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}\ell(A_{i})\right].$ ($\mathcal{P}^{*}$)
Utilizing a distributional strategy grants decision-makers the flexibility to generate any number of viable solutions, circumventing the need to directly model the vast decision space of potential actions. Figure 2(b) provides a graphical illustration of this method, which we denote as generative curation.
@@
Under generative curation, we can show that when desirability is influenced by arbitrary observable and unobservable characteristics, the objective function can be modeled as optimizing a quantitative objective with an added noise term, characterized by a stationary stochastic process. Specifically, we demonstrate that by modeling this stochastic process as Gaussian, the optimization problem becomes a trade-off between quantitative optimality and qualitative diversity, governed by a novel diversity metric derived analytically from the Gaussian process. Our exploration of this diversity metric uncovers that achieving diversity does not merely equate to generating a wide spread of solutions. In fact, even mild assumptions imposed on the Gaussian process can lead to highly non-trivial dynamics in how solution diversity is induced.
To operationalize this framework, we propose two implementation approaches: ($i$) We develop a generative neural network architecture to directly output $\pi$ that can be utilized to sample high-quality solutions in practical decision-making scenarios; ($ii$) We demonstrate that the framework can also be approximated through a sequential optimization approach, where high-quality solutions are iteratively generated. The latter method is particularly adaptable to complex optimization problems, such as routing or assignment tasks, where a distributional formulation of the problem is more difficult. We validate our approach using both synthetic and real-world datasets, demonstrating its effectiveness in enhancing decision-making processes in complex environments.
@@
[…]
(a) Optimal decision-making
(b) Generative Curation
Comparison between classic optimal decision-making and the proposed framework. (a) In classic optimal decision-making settings, optimization models typically focus only on observable quantitative factors with the goal of identifying a single optimal solution. (b) Generative Curation offers multiple thoughtfully curated solutions to human decision-makers, enabling them to choose the most desirable one while actively involving them in the decision-making process.
@@
    Human-Centered Decision Making with Generative Curation: We formalize the concept of human-centered decision making, and considers generating a set of recommendations to maximize human desirability of the action, which could include both quantitative factors and unmeasurable qualitative factors. We introduce the framework of generative curation where we utilize a distributional approach to enable flexible generation of diverse and near-optimal solutions.
@@ 1.1.1 Modeling to Generate Alternatives
”Modeling to generate alternatives” (MGA) is a well-established approach in decision-making that offers policymakers a set of diverse, near-optimal solutions, providing multiple perspectives on a given problem (Brill Jr et al. 1982). Traditionally, MGA research has concentrated on sequential optimization methods (Chang et al. 1982, Greistorfer et al. 2008, Ingmar et al. 2020, Delarue et al. 2023) and integer programming techniques (Trapp and Konrad 2015, Danna et al. 2007, Ahanor et al. 2024) to generate varied yet feasible solutions. In parallel, the recommendation systems literature has investigated integrating diversity objectives into classical algorithms (Puthiya Parambath et al. 2016, Castells et al. 2021). More recently, the Boltzmann distribution has gained popularity for promoting diversity by sampling from probabilistic models (e.g. Mann et al. 2020). Despite the widespread use of predefined metrics such as pairwise distance, entropy, or coverage to quantify diversity, these measures often lack robust justification as proxies for optimal diversity, as Delarue et al. (2023) highlighted. Our research seeks to bridge this gap by proposing a novel diversity metric grounded in a principled modeling framework that captures both observable and unobservable elements of human desirability.
@@ 1.1.3 Human-AI Interactions
More broadly, our work relates to the growing literature on human-AI interactions. Many works have focused on understanding the impact of machine learning (ML) algorithms that provide advice to human decision-makers and explore optimal strategies to combine human and algorithmic decision-making (Grand-Clément and Pauphilet 2024, Orfanoudaki et al. 2022, Chen et al. 2023).
@@ 2.1 Problem Setup
Consider the scenario where a human decision-maker needs to make a decision (or action) $a\in\mathcal{A}\subseteq\mathbb{R}^{d}$ from a large, possibly infinite set of potential actions. The goal is to maximize the human desirability of the actions, $\ell(a)$. Typically, the decision-maker can only formulate a partial part of desirability that depend on some observable features of the actions (and/or the underlying population) $X_{1}(a),\cdots,X_{p}(a)$. There are other important qualitative considerations $V_{1}(a)\sim\mathcal{P}_{V_{1}},\cdots,V_{q}(a)\sim\mathcal{P}_{V_{q}}$ that are unobserved and cannot be easily elicited from data. Without loss of generality, we assume that the underlying true desirability $\ell$ is influenced arbitarily by both components as followed:
@@
where $U$ is a zero-mean stochastic process. If $X_{1}(a),\dots,X_{p}(a)$ are based on observable features, then $Y(a)$ is a deterministic objective function that encodes the quantitative desirability of the action, whereas $U(a)$ retains the variation due to the unobserved qualitative desirability. Figure 3(c)(a) gives an illustrative example of both quantitative and qualitative desirability in an one-dimensional action space.
@@
As detailed in Section 1, the goal of generative curation is to create a probability distribution over the action space $\pi(a):\mathcal{A}\to[0,1]$ that enables the decision-maker to maximize the human desirability of the best action among $m$ sampled actions. We reproduce Problem $\mathcal{P}^{*}$ below:
$\max_{\pi}~{}\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}~{}\ell(A_{i})\right],~{}\text{where}~{}\ell(A_{i})=Y(A_{i})+U(A_{i}).$ ($\mathcal{P}^{*}$)
@@
The value of $m$ is typically limited by practical constraints. For instance, in the police districting problem (Zhu et al. 2020, 2022), the decision-makers were asked to present no more than five districting plans due to time constraints.
@@
This formulation effectively integrates human agency with algorithmic decision-making. The quantitative algorithm generates a policy $\pi$ that samples $m$ actions, from which the human decision-maker selects the one that best aligns with their preferences. We note here that this approach does not assume the decision-maker has precise knowledge of their preference function $\ell(A_{i})$. There is considerable literature that supports individuals often struggle to articulate their own preferences accurately, yet they are significantly better at making comparative judgments between options (e.g., Fischhoff 1991, Tversky and Kahneman 1974, Slovic 1995). This comparative capability is sufficient for selecting the best option among $m$ actions (e.g., through bubble sort as detailed in Astrachan (2003)).

### v2

@@
Operational decisions in healthcare, logistics, and public policy increasingly involve algorithms that recommend candidate solutions, such as treatment plans, delivery routes, or policy options, while leaving the final choice to human decision-makers. For instance, school districts use algorithms to design bus routes, but administrators make the final call given community feedback. In these settings, decision quality depends not on a single algorithmic “optimum”, but on whether the portfolio of recommendations contains at least one option the human ultimately deems desirable. We propose generative curation, a framework that optimally generates recommendation sets when desirability depends on both observable objectives and unobserved qualitative considerations. Instead of a fixed solution, generative curation learns a distribution over solutions designed to maximize the expected desirability of the best option within a manageable portfolio. Our analysis identifies a trade-off between quantitative quality and qualitative diversity, formalized through a novel diversity metric derived from the reformulated objective. We implement the framework using a generative neural network and a sequential optimization method, and show in synthetic and real-world studies that it consistently reduces expected regret compared to existing benchmarks. Our framework provides decision-makers with a principled way to design algorithms that complement, rather than replace, human judgment. By generating portfolios of diverse yet high-quality options, decision-support tools can better accommodate unmodeled factors such as stakeholder preferences, political feasibility, or community acceptance. More broadly, the framework enables organizations to operationalize human-centered decision-making at scale, ensuring that algorithmic recommendations remain useful even when objectives are incomplete or evolving.
@@ 1 Introduction
With the explosion of data availability, human decision-makers faces an ever-growing set of information and decision choices. This phenomenon has spurred significant interest in utilizing data-driven machine learning methods to assist in human decision-making, and there is a quickly expanding body of literature dedicated to developing methods capable of recommending the optimal decision across a variety of scenarios based on a particular objective (e.g., Bertsimas and Kallus 2016, Kube et al. 2019, Bryan and Ventresca 2020).
@@
In this paper, we focus on developing decision recommendations while explicitly accounting for the fact that algorithmic suggestions are subject to human arbitration. We define this category of decision-making algorithms as human-centered decision making. As illustrated in Table 1, this framework contrasts with standard optimal decision-making approaches (e.g., Zhao et al. 2012, Moodie et al. 2012, Zhu et al. 2017, Luedtke and van der Laan 2016, Athey and Wager 2021), as human-centered decision making models problems in which the human decision-maker retains final decision authority.
@@
Algorithm gives
decision input?
No N/A
Human Decision Making
e.g., Doctors, Policymakers
Yes
Optimal Decision Making
e.g., optimal treatment rules
Human-Centered Decision Making
e.g., modeling to generate alternatives, generative curation
Within human-centered decision making, we focus on a common structure: an interactive two-layer process in which operational teams generate a portfolio of $m$ recommended solutions, and a human decision-maker, such as a policymaker, clinician, or executive, selects the most appropriate option based on a combination of quantitative and qualitative criteria. If none of the proposed options are satisfactory, the team is asked to generate a new set of recommendations.
This structure is prevalent in practice because it helps balance two competing needs. On the one hand, decision-makers seek to avoid decision fatigue and minimize cognitive burden by reviewing only a manageable number of options. On the other hand, they also want sufficient diversity in the choices to ensure flexibility and alignment with unmodeled or context-specific considerations. By delegating the generation of candidate plans to the operational team, decision-makers can focus their attention on a smaller, curated set that is easier to evaluate yet broad enough to reflect the complexity of the underlying decision.
For example, in the school bus routing problem studied by Delarue et al. (2024), optimization methods are used to generate feasible transportation plans, but the final selection is made by school administrators, who weigh additional factors such as parent feedback, implementation feasibility, and stakeholder coordination. Similar decision workflows are common in domains such as urban planning (Lin et al. 2022), clinical guideline development (Wu et al. 2024), and emergency response planning (Bertsimas et al. 2022, Zhu et al. 2022), where algorithmic support is paired with human judgment to navigate multidimensional trade-offs.
To support such decision structures, some approaches attempt to elicit and model multiple human preference dimensions explicitly, presenting diverse solutions along a Pareto frontier (Masin and Bukchin 2008, Lin et al. 2022, Wu et al. 2024). While these methods can be valuable when the dimensions of human preferences are well understood, they tend to be less effective when qualitative dimensions are ambiguous or difficult to articulate in advance. In these cases, the true preference function of humans, $\ell(\cdot)$, is partially unknowable. For instance, during the COVID-19 pandemic, governments adopted policies shaped not just by epidemiological metrics but also by unquantifiable concerns such as public and political feasibility (Bertsimas et al. 2022). In police districting, plans optimized for workload balance (Larson 1974, Zhu et al. 2020) have been rejected due to practical concerns like highway access and neighborhood integrity that are difficult, if not impossible to quantify (Shirabe 2009, Gardner et al. 2014). In such settings, the existing practices often resort to heuristic rules or ad hoc constraints to produce a range of solutions that are near-optimal in the quantifiable parts of the objective, and allow decision-makers to choose from such set (Chang et al. 1982, Greistorfer et al. 2008).
In this paper, we propose a new framework to formalize how to optimally support these two-layer decision-making structures when underlying qualitative factors are unknowable or difficult to model. Our starting point is the observation that the goal of a recommendation system in a human-centered decision-making context is not to identify a single best solution, but rather to facilitate human decisions that are ultimately deemed most desirable by the decision-maker. Consequently, the recommendation team tries to solve the objective of finding a set of $m$ candidate actions $A_{1},\cdots,A_{m}\in\mathcal{A}$ such that, once the human evaluates them, the most desirable option in the set maximizes the human-assessed desirability function $\ell(\cdot)$:
$\max_{A_{1},\cdots,A_{m}\in\mathcal{A}}~\max_{1\leq i\leq m}\ell(A_{i}).$ (1)
If the human desirability function $\ell(\cdot)$ were fully known, this objective would reduce to a standard one-layer optimization problem: one would simply identify the action $A$ that maximizes $\ell(A)$. However, since $\ell(\cdot)$ is partially unobservable, Problem 1 diverges meaningfully from standard optimal decision-making formulations. Any fixed set of $m$ recommendations carries the risk of failing to meet the decision-maker’s expectations. When that occurs, the human decision-maker typically requests a new batch of alternatives, prompting the team to generate additional plans.
@@
Workflow of the proposed generative curation.
This iterative interaction between human and algorithm highlights the need for a mechanism that can reliably produce new sets of candidate actions until a satisfactory option emerges. This insight motivates our proposal to construct an optimal generative distribution $\pi$ that can be repeatedly sampled to produce diverse and high-quality recommendation sets. Such a generative structure equips the recommendation team with a tool that can quickly generate new solutions in case the old solutions were not satisfactory. Figure 1 illustrates this interactive workflow. Formally, our goal is therefore to maximize the expected human desirability of the best action within a set of sampled recommendations sampled from $\pi$:
$\max_{\pi}~\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}\ell(A_{i})\right].$ (2)
where the expectation is taken over both the randomness in the generative distribution $\pi$ and the uncertainty in the unobserved portion of the human desirability function $\ell(\cdot)$. This formulation provides a principled foundation for designing recommendation systems that support human-centered decision-making under partial preference observability. We denote this formulation as generative curation.
@@
Under generative curation, we show that when desirability is influenced by both observable and unobservable characteristics, the objective function can be modeled as optimizing a known quantitative objective augmented by a noise term, characterized by a stationary stochastic process. In particular, by modeling this stochastic process as Gaussian, the resulting optimization problem reveals an inherent trade-off between quantitative optimality and qualitative diversity. This trade-off is governed by a novel diversity metric, which we derive analytically from the properties of the Gaussian process.
Our analysis of this metric uncovers a key insight: achieving qualitative diversity is not simply a matter of generating a wide spread of solutions. Even under mild assumptions about the Gaussian process, the relationship between diversity and desirability can exhibit complex and nontrivial behavior. These dynamics highlight the importance of structuring diversity carefully, so that it reflects robustness to unknown preferences rather than arbitrary variation.
To operationalize the generative curation framework, we propose two implementation approaches. First, we develop a generative neural network architecture that directly outputs the distribution $\pi$, enabling efficient sampling of high-quality and diverse recommendations in practical decision-making settings. Second, we present a sequential optimization method that approximates the generative objective by iteratively generating desirable alternatives. This approach is particularly well suited for structured combinatorial problems, such as routing and assignment, where directly modeling a distribution $\pi$ may be intractable.
We validate both implementations on a range of synthetic and real-world datasets, demonstrating that generative curation consistently produces recommendation sets that improve alignment with human preferences in settings where qualitative factors are difficult to model explicitly. Our results show that the framework is broadly applicable and effective in enhancing decision quality in complex, high-stakes environments.
@@
    Human-Centered Decision Making with Generative Curation: We formalize the concept of human-centered decision making, and considers generating a set of recommendations to maximize human desirability of the action, which could include both quantitative factors and unmeasurable qualitative factors. We introduce the framework of generative curation to optimally recommend human decision-makers in a generative manner.
@@ 1.1.1 Modeling to Generate Alternatives
“Modeling to generate alternatives” (MGA) is a well-established approach in decision-making that offers policymakers a set of diverse, near-optimal solutions, providing multiple perspectives on a given problem (Brill Jr et al. 1982). Traditionally, MGA research has concentrated on sequential optimization methods (Chang et al. 1982, Greistorfer et al. 2008, Ingmar et al. 2020, Delarue et al. 2023) and integer programming techniques (Trapp and Konrad 2015, Danna et al. 2007, Ahanor et al. 2024, Wang et al. 2024) to generate varied yet feasible solutions. In parallel, the recommendation systems literature has investigated integrating diversity objectives into classical algorithms (Puthiya Parambath et al. 2016, Castells et al. 2021). More recently, the Boltzmann distribution has gained popularity for promoting diversity by sampling from probabilistic models (e.g., Mann et al. 2020). Despite the widespread use of predefined metrics such as pairwise distance, entropy, or coverage to quantify diversity, these measures often lack robust justification as proxies for optimal diversity, as Delarue et al. (2023) highlighted. Our research seeks to bridge this gap by proposing a novel diversity metric grounded in a principled modeling framework that captures both observable and unobservable elements of human desirability.
@@ 1.1.3 Human-AI Interactions
More broadly, our work relates to the growing literature on human-AI interactions. Many works have focused on understanding the impact of machine learning algorithms that provide advice to human decision-makers and explore optimal strategies to combine human and algorithmic decision-making (Grand-Clément and Pauphilet 2024, Orfanoudaki et al. 2022, Chen et al. 2023, Te’eni et al. 2023).
@@ 2.1 Problem Setup
Consider the scenario where a human decision-maker needs to make a decision (or action) $a\in\mathcal{A}\subseteq\mathbb{R}^{d}$ from a large, possibly infinite set of potential actions. The goal is to maximize the human desirability of the actions, $\ell(a)$. Typically, the decision-maker can only formulate a partial part of desirability that depend on some observable, deterministic features of the actions (and/or the underlying population) $X_{1}(a),\cdots,X_{p}(a)$. For every action $a$, there are other important qualitative considerations (e.g., implementation feasibility) $V_{1}(a),\cdots,V_{q}(a)$ that cannot be easily elicited and are unknown apriori. In particular, we assume that $(V_{1}(a),\cdots,V_{q}(a))$ follows an unknown joint prior distribution $\mathcal{P}_{V}(a)$. Without loss of generality, the underlying true desirability $\ell$ is influenced arbitrarily by both components as followed:
@@
where $U$ is a zero-mean stochastic process (indexed by $a$) and $Y(a)$ is a function that only depends on $X_{1}(a),\cdots,X_{p}(a)$. Therefore, $Y(a)$ is a deterministic objective function that encodes the quantitative desirability of the action, whereas $U(a)$ retains the variation due to the unobserved qualitative desirability. Figure 2(c)(a) gives an illustrative example of both quantitative and qualitative desirability in an one-dimensional action space.
@@
As detailed in Section 1, the goal of generative curation is to create a probability distribution over the action space $\pi(a):\mathcal{A}\to[0,1]$ that enables the decision-maker to maximize the human desirability of the best action among $m$ sampled actions. We reproduce Problem 2 below:
$\max_{\pi}~\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}~\ell(A_{i})\coloneqq Y(A_{i})+U(A_{i})\right],$
@@
Under our setup, the expectation is taken over both the randomness in sampling actions $a$ over $\pi$ and also the uncertainty in the qualitative desirability $V_{1}(a),\cdots,V_{q}(a)$. The value of $m$ is typically limited by practical constraints. For instance, in the police districting problem (Zhu et al. 2020, 2022), the decision-makers were asked to present no more than five districting plans due to time constraints.
@@
This formulation effectively integrates human agency with algorithmic decision-making. The quantitative algorithm generates a policy $\pi$ that samples $m$ actions $A_{1},\cdots,A_{m}$, from which the human decision-maker selects the one that best aligns with their preferences. We note here that this approach does not assume the decision-maker has precise knowledge of their true preference function $\ell(A_{i})$, but only assumes that the decision-maker can compare different options and choose the one that best aligns with their true preference function $\ell(A_{i})$. There is considerable literature that supports individuals often struggle to articulate their own preferences accurately, yet they are significantly better at making comparative judgments between options (e.g., Fischhoff 1991, Tversky and Kahneman 1974, Slovic 1995). This comparative capability is sufficient for selecting the best option among $m$ actions (e.g., through bubble sort as detailed in Astrachan (2003)).
2.2 A Tight Reformulation with Gaussian Processes
To solve Problem 2, it is necessary to impose certain assumptions on the structure of the zero-mean stochastic process $U(A_{i})$. Specifically, we assume that this process can be modeled as a zero-mean, stationary Gaussian process (GP) with a covariance function $k(a,a^{\prime})$: {assumption} The qualitative desirability $U(a)$ can be modeled by a stationary GP with a mean of zero and a covariance function $k(a,a^{\prime})$. Compared with the general formulation of $U(a)$ being a zero-mean stochastic process, Assumption 2.2 is equivalent to the addition of two conditions (see e.g., Edition et al. 2002): ($i$) For any set of any number of actions $a_{1},\cdots,a_{n}$, $U(a_{1}),\cdots,U(a_{n})$ follows a multivariate normal distribution. ($ii$) $U(a)$ is second-order stationary – i.e., $\operatorname{Cov}(U(a_{1}),U(a_{2}))$ only depends on the distance $\|a_{1}-a_{2}\|$.
Condition ($i$) ensures that the qualitative desirability across actions is normally distributed, implying that extreme outliers in desirability are unlikely. There is significant prior work that utilize this normality assumption to model human preferences (Chu and Ghahramani 2005, González et al. 2017, Bıyık et al. 2020).
Condition ($ii$), second-order stationarity, reflects a minimal prior information principle. In the absence of detailed knowledge about the qualitative desirability function, it is reasonable to assume that correlations between actions depend only on their relative proximity, rather than their specific identities. While this assumption is useful when prior knowledge is scarce, it can limit the model’s ability to capture complex, non-uniform correlations. If prior information on non-uniform correlations is available, the decision space can be divided into regions where stationarity holds locally, allowing the model to be applied separately in each region.